OpenStudy (anj123):

What is the derivative of 2x^(-1/2) +x^(3/2) ? Please show steps...

OpenStudy (agent0smith):

Just use the power rule \[\large \frac{ d }{dx }\left( x^n \right) = n*x^{n-1}\]

OpenStudy (anj123):

So, the answer is -x^(-3/2)+3/2(x^(1/2))

OpenStudy (anj123):

the answer is -x^(-3/2) + (3/4)x^(-1/2)

OpenStudy (anj123):

is the second answer correct?

OpenStudy (agent0smith):

Neither look correct

OpenStudy (anj123):

okay so the first part..... 2x^(-1/2) the derivative of that is -x^(-3/2) right?

OpenStudy (agent0smith):

Right

OpenStudy (anj123):

the second part is (3/2)x^1/2 the derivative of the second part is (3/4)x^(-1/2) correct?

OpenStudy (agent0smith):

the second part is x^(3/2) Find the deriv. of that.